# AI-Driven Control of Chaos: A Transformer-Based Approach for Dynamical Systems

### v1

@@ Abstract
Chaotic behavior in dynamical systems poses a significant challenge in trajectory control, traditionally relying on computationally intensive physical models. We present a machine learning-based algorithm to compute the minimum control bounds required to confine particles within a region indefinitely, using only samples of orbits that iterate within the region before diverging. This model-free approach achieves high accuracy, with a mean squared error of $2.88\times 10^{-4}$ and computation times in the range of seconds. The results highlight its efficiency and potential for real-time control of chaotic systems.
pacs:
$05.45.-a$, $05.45.Gg$, $07.05.Mh$
Chaotic transients, characterized by their unpredictable behavior, present challenges to trajectory control and predictability in dynamical systems, particularly in real-time or high-dimensional contexts. Traditional control methods depend on computationally expensive physical models, limiting their practicality. This study introduces a machine learning framework utilizing a transformer architecture to estimate control bounds directly from data, bypassing explicit models. The approach is robust to noise and adaptable to varying conditions, offering scalable solutions for managing chaotic transients in complex systems.
I Introduction
@@
The study of chaos and the development of methods to predict and control chaotic systems have a rich history, with significant progress made over the past few decades. Early approaches focused on understanding the mathematical foundations of chaos, leading to the identification of key properties such as Lyapunov exponents and strange attractors, which describe the system’s sensitivity to initial conditions and the structure of chaotic trajectories  peitgen1992chaos . These foundational concepts paved the way for the development of control strategies aimed at stabilizing chaotic systems Ott1990 ; Shinbrot1990 ; Braiman1991 ; Pyragas1992 ; Shinbrot1993 ; Grebogi1997 .
Building on these early successes, researchers have developed various methods to control chaotic systems, with significant focus on managing chaotic transients. These transients are periods of chaotic motion that occur temporarily before a system stabilizes or escapes to another attractor.
@@
Managing chaotic transients is essential in systems where maintaining a chaotic regime offers significant benefits. For example, in mechanical systems, chaotic vibrations can improve the efficiency of energy harvesters by broadening the range of resonance frequencies and increasing energy extraction litak2015regular . Similarly, in ecological models, chaotic dynamics can help prevent population collapses and maintain ecosystem stability capeans2019new . Among the most effective techniques developed to achieve these goals is partial control sabuco2012dynamics ; capeans2017partially , which utilizes tools such as safety functions and safe sets capeans2019new .
The safety function is a powerful tool for controlling systems with chaotic transients governed by difference equations. It quantifies the minimum control required to confine chaotic trajectories within an arbitrary region of phase space for a given number of iterations capeans2022controlling ; Capeans2024 .
@@
While this approach represents a significant advancement in the control of chaotic systems, the computation of safety functions remains highly resource intensive. This challenge becomes particularly evident in systems with a high number of dimensions or variables, as well as in scenarios requiring real-time decision making. As a result, this computational bottleneck limits the broader adoption of safety functions in fields that demand rapid and efficient solutions.
@@
In this work, we aim to address these computational challenges by incorporating machine learning techniques to estimate safety functions. Using data collected from various chaotic systems, supervised training of the models provides an efficient approximation of the safety function. Samples for the training consists of orbits of variable length escaping a predefined region. The models learn dynamical patterns that generalize across different systems. The primary goal is to reduce the computational burden of safety function calculations while maintaining accuracy, making the approach feasible for real-time and large scale applications.
@@
Our methodology involves collecting datasets from systems with chaotic transients, training models with varying levels of available information, and evaluating their performance. By focusing on generalizability, we aim to create predictive models of safety functions from time series data for a wide range of systems without requiring fine tuning. This approach not only improves efficiency but also expands the practical utility of safety functions in controlling chaotic transients.
@@
The structure of this paper is as follows. Section II reviews the current state of the art in the computation of safety functions and discusses the role of machine learning in addressing these challenges. Section III details the data collection, model design, and training processes. In Sec. IV, we evaluate the predictive models, examining their performance and efficiency across different scenarios. Finally, Sec. V summarizes our findings, discusses their implications, and outlines potential avenues for future work.
II Theoretical background
Transient chaos refers to the temporary chaotic behavior exhibited by a system before it stabilizes or escapes to an specific region. This phenomenon presents unique challenges when combined with the intrinsic instability of chaotic systems and the unavoidable presence of noise, which exists in any real experimental setup or dataset.
@@ II Theoretical background
Maintaining a small control under these conditions may seem unrealistic. However, it was discovered that a region $Q$, which contains a chaotic saddle, also includes a special subset where chaotic trajectories can be sustained with a minimal amount of control capeans2018partial . This subset, known as the safe set, depends on both the noise strength affecting the system and the available control.
The partial control method focuses on keeping chaotic trajectories confined within the safe set using minimal control even when noise is present. Surprisingly, partial control can find safe sets where the magnitude of control needed to prevent the chaotic orbit from escaping is smaller than the noise itself.
@@
A recent approach enhances the partial control method by introducing a safety function, which automatically computes the minimum control required for each point in the region $Q$ to remain within the desired bounds. This method has been applied to well known chaotic systems such as the Hénon and Lozi maps, demonstrating that chaotic dynamics can be controlled with minimal intervention sabuco2012dynamics ; capeans2022controlling .
Given a dynamical system, and a region $Q$ where we aim to confine the system orbits forever, we define the safety function for that region through the recursive equation:
$U_{k+1}=f(\xi,U_{k})$ (1)
Starting with $U_{0}=0$, where $U_{k}$ represents the safety function for the region $Q$, and thus the minimum control required for each initial condition in $Q$ to keep its orbit within the chaotic transient for $k$ iterations. Since our goal is to maintain the trajectory in $Q$ indefinitely, we need to find $U_{\infty}$, which is obtained when this recursive equation converges. The function $U_{\infty}$ is the function targeted for optimization in this paper, it depends on the upper bound of noise $\xi$ present in the system. A more detailed explanation of the computation and derivation of this function is provided in Appendix A.
@@
This map exhibits transient chaos within the interval $Q=[0,1]$. In Fig. 1(a), we illustrate how an orbit undergoes chaotic motion until it eventually exits $Q$. Figures 1(b-c) depict the same safety function associated with this map under noise $\xi=0.05$. To identify a safe set, we extract the set of initial conditions for which the safety function is below a predefined value $u$ (represented by a red line) such that in each iteration of the orbit $u_{n}\leq u$, we arbitrarily set $u=0.035$ Fig. 1(b) and $u=0.07$ in Fig. 1(c). The resulting safe set, highlighted in yellow, confines the trajectory indefinitely within $Q$.
Figures 1(d-e) demonstrate how the orbit iterates within the map, bouncing within the regions of the safe set indefinitely, in contrast to the case shown in Fig. 1(a) where no control is applied and the orbit diverges. Finally, Figs. 1(f-g) show the value of the control applied during the first $100$ iterations. This highlights that the control method consistently remains within the preselected maximum $u$ value, mantaining the trajectory in the region of the space where transient chaos is present.
@@
Machine learning offers a viable solution to these challenges by approximating the convergence behavior. This approach has been successfully employed in various fields of physics to optimize complex processes. For instance, neural networks have been used to represent quantum wavefunctions, accelerating the estimation of ground states in many body quantum systems carleo2017solving . Convolutional neural networks have been applied to characterize the complexity of the basins of attraction in chaotic systems, significantly outperforming classical algorithms in processing speed Valle2022 ; Valle2024 . Deep reinforcement learning has also been leveraged to tackle complex control problems in plasma physics, enabling the control of magnetic fields and stabilization of plasma in nuclear fusion reactors degrave2022magnetic .
These applications show how machine learning can reduce the computational demands of iterative estimation processes and optimize control tasks in high-dimensional physical systems, highlighting the potential of similar methods for convergence estimation in dynamical systems like the one under study.
A major advancement in the field has been the introduction of transformer models, as described by Vaswani et al. vaswani2017attention . Transformers revolutionized natural language processing by replacing recurrent and convolutional layers with a self-attention mechanism, capturing dependencies in sequences regardless of distance. This innovation improved parallelizability and reduced training times compared to earlier architectures like Long Short-Term Memory networks  hochreiter1997long and Convolutional Neural Networks lecun1998gradient .
Figure 1: Illustration of transient chaos and the application of the safety function for the tent map with noise $\bm{\xi=0.05}$ in the interval $\bm{Q=[0,1]}$. Panel $(a)$ shows a non controlled chaotic trajectory of the noisy map, eventually escaping region $Q$. Panels $(b)$ and $(c)$ display the safety function $U_{\infty}$ for the map in $(a)$, with maximum control values arbitrarily set to $u=0.035$ and $u=0.07$, respectively (red line). The yellow regions represent the safe set, corresponding to areas where the safety function lies below the maximum control value $u$, allowing trajectories to remain indefinitely when control is applied. Panels $(d)$ and $(e)$ show two trajectories confined to the safe set, with control $u_{n}\leq u$ applied at each iteration, contrasting the escape seen in $(a)$. Panels $(f)$ and $(g)$ illustrate the control values applied during the first $100$ iterations, confirming that they remain consistently below the selected maximum $u$. This demonstrates the effectiveness of the safety function in maintaining trajectories within region $Q$.
@@
III Methodology
This work has been fully implemented in Python, with all scripts and data publicly available in a GitHub repository RedLynx96_Estimation_of_safety_functions .
To generate the datasets, we developed a Python script that creates one-dimensional pseudorandom physical models. For these systems, we collect samples of orbits that iterate within a predefined region $Q$ until they escape, along with the corresponding safety function for this region. Additional details about the generation of pseudorandom functions, sampled orbits, and safety functions can be found in Appendix B.
@@ III Methodology
The architecture comprises two attention blocks, followed by a convolution block, a pooling block, and a final output block of fully connected layers. This configuration was chosen as it empirically provided the best performance during preliminary experiments. The two attention blocks effectively capture complex, long range dependencies in the input time series, while the subsequent convolution block complements this by extracting localized patterns. This balance allows the model to generalize well and accurately predict the safety function.
Figure 2: Architecture of the neural network used in this research. The model includes two input branches: Input 1, representing the time series of sampled orbits with array length Mean_ext, and Input 2, the initial condition for which the control value is calculated. Input 2 is repeated and merged with Input 1 before being processed through two transformer blocks, a convolutional block, a pooling block, and a final output block. Each layer’s output shape is shown below its name. Key components include Multi Head Attention layers to capture dependencies in the time series, Conv1D layers for local pattern extraction, and pooling layers to reduce dimensionality. Dense layers with Dropout finalize the prediction of the minimum control value required to keep the initial condition within the safe set indefinitely.
@@
All network weights and biases were randomly initialized and optimized using TensorFlow’s Adam optimizer  kingma2014adam , configured with the following parameters: learning rate $=lr_{0}=1.5\cdot 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, $\text{decay}=0$, and $amsgrad=\text{true}$. A warmup phase was introduced at the start of training to adapt the learning process to the transformer architecture. The tuning of the learning rate is described below:
$lr_{t}=\begin{cases}lr_{0}\cdot\frac{t}{W_{e}}&\text{for }t\leq W_{e}\\ lr_{0}\cdot d^{\frac{t-W_{e}}{T-W_{e}}}&\text{for }t>W_{e}\end{cases}$ (3)
Here $lr_{t}$ denotes the learning rate at epoch $t\geq 1$, starting at an initial value of $lr_{0}$. The learning rate increases linearly during the $W_{e}=6$ warmup epochs, followed by a decay phase controlled by a decay factor $d=0.2$. This warmup phase is an effective step for the training of transformer models, as it helps stabilize optimization by gradually increasing the learning rate during the initial training phase, preventing issues with large gradients  vaswani2017attention .
@@
$\text{Wmse}=\frac{1}{M}\sum_{i=1}^{M}w_{i}\cdot(y_{t,i}-y_{p,i})^{2}$ (4)
@@
$w_{i}=\begin{cases}10&\text{if }(y_{t,i}-y_{p,i})>0\\ 1&\text{if }(y_{t,i}-y_{p,i})\leq 0\end{cases}$ (5)
This weighting approach introduces a bias for the overestimation of the safety function. Underestimation can lead to an orbit escaping early the region $Q$. Since the safety function defines the maximum control bound needed to confine the orbit, underestimations pose a critical risk. To mitigate this, the training process imposes a significantly higher penalty on underestimations, prompting the model to prioritize larger predictions. This conservative choice for the estimation lead to a satisfactory trade-off in performance as we will show in the next section.
IV Results
As detailed in Sec III, the transformer-based model illustrated in Fig. 2 was trained on different amounts of sampled data. We trained $6$ models, each with progressively more data than the last. This approach simulates real world scenarios where data availability may be limited.

### v2

@@ Abstract
Chaotic behavior in dynamical systems poses a significant challenge in trajectory control, traditionally relying on computationally intensive physical models. We present a machine learning-based algorithm to compute the minimum control bounds required to confine particles within a region indefinitely, using only the first iterations of diverging orbits as required information of the system. This model-free approach achieves high accuracy, with a mean squared error of $2.88\times 10^{-4}$ and computation times in the range of seconds. The results highlight its efficiency and potential for real-time control of chaotic systems.
keywords:
Chaos Control , Partial Control , Machine Learning , Neural Networks , Nonlinear Dynamics
1 Introduction
@@
The study of chaos and the development of methods to predict and control chaotic systems have a rich history, with significant progress made over the past few decades. Early approaches focused on understanding the mathematical foundations of chaos, leading to the identification of key properties such as Lyapunov exponents and strange attractors, which describe the system’s sensitivity to initial conditions and the structure of chaotic trajectories [1]. These foundational concepts paved the way for the development of control strategies aimed at stabilizing chaotic systems [2, 3, 4, 5, 6, 7].
@@
Building on these early successes, researchers have developed various methods to control chaotic systems, with significant focus on managing chaotic transients. These transients are periods of chaotic motion that occur temporarily before a system stabilizes or escapes to another attractor. Managing chaotic transients is essential in systems where maintaining a chaotic regime offers significant benefits. For example, in mechanical systems, chaotic vibrations can improve the efficiency of energy harvesters by broadening the range of resonance frequencies and increasing energy extraction [8]. Similarly, in ecological models, chaotic dynamics can help prevent population collapses and maintain ecosystem stability [9]. Among the most effective techniques developed to achieve these goals is partial control [10, 11], with safety functions and safe sets as their core elements [12].
@@
The safety function is a powerful tool for controlling systems with chaotic transients governed by difference equations. It quantifies the minimum control required to confine chaotic trajectories within an arbitrary region of phase space for a given number of iterations [13, 14]. While this approach represents a significant advancement in the control of chaotic systems, the computation of safety functions remains highly resource intensive. This challenge becomes particularly evident in systems with a high number of dimensions, as well as in scenarios requiring real-time decision making. As a result, this computational bottleneck limits the broader adoption of safety functions in fields that demand rapid and efficient solutions.
@@
In this work, we aim to address these challenges by incorporating machine learning techniques to estimate safety functions. The primary goal is to reduce the computational burden of safety function calculations while maintaining accuracy, making the approach feasible for real-time and large scale applications.
@@
Our methodology involves collecting datasets from unstable systems with chaotic transients, training models with varying levels of available information, and evaluating their performance. By focusing on generalization, we create predictive models of safety functions for a wide range of systems without requiring fine tuning. Once the models have been trained, the control of a system only requires samples of diverging trajectories of the uncontrolled system and the actual state. This approach not only improves efficiency but also expands the practical utility of safety functions in controlling chaotic transients.
@@
The structure of this paper is as follows. Section 2 reviews the current state of the art in the computation of safety functions and discusses the role of machine learning in addressing these challenges. Section 3 details the data collection, model design, and training processes. In Sec. 4, we evaluate the predictive models, examining their performance and efficiency across different scenarios. Finally, Sec. 5 summarizes our findings, discusses their implications, and outlines potential avenues for future work.
2 Theoretical background
@@
Transient chaos refers to the temporary chaotic behavior exhibited by a system before it stabilizes or escapes to a specific region. This phenomenon presents unique challenges when combined with the intrinsic instability of chaotic systems and the unavoidable presence of noise, which exists in any real experimental setup or dataset. Maintaining a small control under these conditions may seem unrealistic. However, it was discovered that a region $Q$, which contains a chaotic saddle, also includes a special subset where chaotic trajectories can be sustained with a minimal amount of control [10, 15]. This subset, known as the safe set, depends on both the noise strength affecting the system and the available control.
@@
The partial control method focuses on keeping chaotic trajectories confined within the safe set using minimal control even when noise is present. Surprisingly, partial control can find safe sets where the magnitude of control needed to prevent the chaotic orbit from escaping is smaller than the noise itself. A recent approach [12] enhances the partial control method by introducing a safety function, which automatically computes the minimum control required for each point in the region $Q$ to remain within the desired bounds. This method has been applied to well known chaotic systems such as the Hénon and Lozi maps, demonstrating that chaotic dynamics can be controlled with minimal intervention [13].
Consider the noisy, partially controlled discrete map:
$q_{n+1}=f(q_{n},\xi_{n})+u_{n},\qquad|\xi_{n}|\leq\xi,$ (1)
where $f$ is the deterministic map, $\xi_{n}$ is an additive disturbance bounded by the noise amplitude $\xi$ and $u_{n}$ is the control effort applied at each iteration.
Without control most trajectories eventually leave the interval $Q$ displaying transient chaos. To quantify how much control is required at every point $q_{i}\in Q$, we introduce the sequence of functions $\{U_{k}\}$ generated by the recursion
$U_{0}(q_{i})=0,\qquad\forall\,q_{i}\in Q,$ (2)
$U_{k+1}(q_{i})=g(U_{k},q_{i},\xi).$
Here, the operator $g$ implements the classical safety function algorithm, as detailed in A, which is the target of optimization in this paper. For a finite horizon $k$, the value $U_{k}(q_{i})$ represents the minimum control upper bound required to ensure that the trajectory starting at $q_{i}$ remains within the set $Q$ for at least $k$ iterations. The corresponding sequence of $k$ controls applied satisfies that independently of the realization of the noise:
$\max(|u_{1}|,|u_{2}|,\dots,|u_{k}|)\leq U_{k}(q_{i}).$ (3)
To ensure that the system’s trajectory remains indefinitely within the set $Q$, we consider the limit of the sequence $\{U_{k}\}$ as $k\to\infty$. Fortunately, we do not need to compute infinite functions, as the presence of noise induces convergence of this sequence after a finite number of steps. That is, there exists some finite value of $k$ such that $U_{\infty}=U_{k+1}=U_{k}$. This function $U_{\infty}$ is called the safety function. For each $q_{i}$ the value $U_{\infty}(q_{i})$ gives the smallest control upper bound that guarantees confinement in $Q$ for an arbitrarily long time.
Once the safety function $U_{\infty}$ has been determined, the control of a trajectory follows immediately. With the only knowledge of the actual perturbed state $f(q_{i},\xi_{n})$, we have to evaluate every admissible controlled image:
$q_{j}=f\bigl{(}q_{i},\xi_{n}\bigr{)}+\;u_{n},\qquad q_{j}\in Q,$
and choose, among all feasible controls $u_{n}$, the one that minimizes the worst-case upper bound:
$\underset{q_{j}\in Q}{\min}\;\Bigl{(}\max\bigl{(}|u_{n}|,\,U_{\infty}(q_{j})\bigr{)}\Bigr{)}.$ (4)
Applying this rule at every iteration guarantees that the entire trajectory starting at $q_{i}$ remains in $Q$ while satisfying the bound $\lvert u_{n}\rvert\leq U_{\infty}(q_{i})$.
Moreover, if we focus only on those states that can be stabilized with an amplitude not exceeding a prescribed value $u$, we define the safe set $\mathcal{S}(u)$ as:
$\mathcal{S}(u)=\bigl{\{}\,q\in Q:U_{\infty}(q)\leq u\,\bigr{\}}.$ (5)
Thus, any orbit starting in $\mathcal{S}(u)$ can be maintained within $\mathcal{S}(u)$ indefinitely by applying a control input not exceeding $u$ at each iteration of the map, as defined in Eq. (1).
@@
Where $x_{n}$ is the state and $u_{n}\leq u$ is the control applied at each iteration. This map exhibits transient chaos within the interval $Q=[0,1]$. In Fig. 1(a), we illustrate how an orbit undergoes chaotic motion until it eventually exits $Q$.
Figure 1: Illustration of transient chaos and the application of the safety function for the tent map with noise $\bm{\xi=0.05}$ in the interval $\bm{Q=[0,1]}$. Panel $(a)$ shows a non controlled chaotic trajectory of the noisy map, eventually escaping the region $Q$. Panels $(b)$ and $(c)$ display the safety function $U_{\infty}$ for the map in $(a)$, with maximum control values arbitrarily set to $u=0.035$ and $u=0.07$, respectively (red line). The yellow regions represent the safe set, corresponding to areas where the safety function lies below the maximum control value $u$, allowing trajectories to remain indefinitely when control is applied. Panels $(d)$ and $(e)$ show two trajectories confined to the safe set, with control $u_{n}\leq u$ applied at each iteration, contrasting the escape seen in $(a)$. Panels $(f)$ and $(g)$ illustrate the control values applied during the first $100$ iterations, confirming that they remain consistently below the selected maximum $u$. This demonstrates the effectiveness of the safety function in maintaining trajectories within region $Q$.
Figures 1(b-c) depict the same safety function associated with this map using a noise $\xi=0.05$. To identify a safe set $\mathcal{S}(u)$, we extract the set of initial conditions for which the safety function is below a predefined value $u$ (represented by a red line), we arbitrarily set $u=0.035$ in Fig. 1(b) and $u=0.07$ in Fig. 1(c). The resulting safe set, highlighted in yellow, is where the trajectory is confined indefinitely.
Figures 1(d-e) demonstrate how the orbit iterates within the map, bouncing within the regions of the safe set indefinitely, in contrast to the case shown in Fig. 1(a) where no control is applied and the orbit diverges. Finally, Figs. 1(f-g) show the value of the control applied during the first $n=100$ iterations.
@@
Machine learning offers a viable solution to these challenges by approximating the convergence behavior. This approach has been successfully employed in various fields of physics to optimize complex processes. For instance, neural networks have been used to represent quantum wavefunctions, accelerating the estimation of ground states in many body quantum systems [16]. Convolutional neural networks have been applied to characterize the complexity of the basins of attraction in chaotic systems, significantly outperforming classical algorithms in processing speed [17, 18]. Deep reinforcement learning has also been leveraged to tackle complex control problems in plasma physics, enabling the control of magnetic fields and stabilization of plasma in nuclear fusion reactors [19].
These applications show how machine learning can reduce the computational demands of iterative estimation processes and optimize control tasks in high-dimensional physical systems, highlighting the potential of similar methods for convergence estimation in dynamical systems like the one under study. A major advancement in the field has been the introduction of transformer models, as described by Vaswani et al. [20]. Transformers revolutionized natural language processing by replacing recurrent and convolutional layers with a self-attention mechanism, capturing dependencies in sequences regardless of distance. This innovation improved parallelizability and reduced training times compared to earlier architectures like Long Short-Term Memory networks [21] and Convolutional Neural Networks [22].
@@
3 Methodology
This work has been fully implemented in Python, with all scripts and data publicly available in a GitHub repository [23].
To generate the datasets, we developed a Python script that creates one-dimensional pseudorandom physical models. For these systems, we collect samples of orbits that iterate within a predefined region $Q$ until they escape, along with the reference safety function for this region. Additional details about the generation of pseudorandom functions, sampled orbits, and safety functions can be found in B.
@@ 3 Methodology
The architecture comprises a stack of two transformer blocks, followed by a convolution block, a pooling block, and a final output block of fully connected layers. This configuration was chosen as it empirically provided the best performance during preliminary experiments. The two transformer blocks effectively capture complex, long range dependencies in the input time series, while the subsequent convolution block complements this by extracting localized patterns. This balance allows the model to generalize well and accurately predict the safety function.
@@
Figure 2: Architecture of the neural network used in this research. The model includes two input branches: Input 1, representing the time series of sampled orbits with array length Mean_ext, and Input 2, the initial condition for which the control value is calculated. Input 2 is repeated and merged with Input 1 before being processed through a stack of two transformer blocks, a convolutional block, a pooling block, and a final output block. Each layer’s output shape is shown below its name. Key components include Multi Head Attention layers to capture dependencies in the time series, Conv1D layers for local pattern extraction, and pooling layers to reduce dimensionality. Dense layers with Dropout finalize the prediction of the minimum control value required to keep the initial condition within the safe set indefinitely.
@@
All network weights and biases were randomly initialized and optimized using TensorFlow’s Adam optimizer [24], configured with the following parameters: learning rate $=lr_{0}=1.5\cdot 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, $\text{decay}=0$, and $amsgrad=\text{true}$. A warmup phase was introduced at the start of training to adapt the learning process to the transformer architecture. The tuning of the learning rate is described below:
$lr_{t}=\begin{cases}lr_{0}\cdot\frac{t}{W_{e}}&\text{for }t\leq W_{e}\\ lr_{0}\cdot d^{\frac{t-W_{e}}{T-W_{e}}}&\text{for }t>W_{e}.\end{cases}$ (7)
Here $lr_{t}$ denotes the learning rate at epoch $t\geq 1$, starting at an initial value of $lr_{0}$. The learning rate increases linearly during the $W_{e}=6$ warmup epochs, followed by a decay phase controlled by a decay factor $d=0.2$. This warmup phase is an effective step for the training of transformer models, as it helps stabilize optimization by gradually increasing the learning rate during the initial training phase, preventing issues with large gradients [20].
@@
$\text{Wmse}=\frac{1}{M}\sum_{i=1}^{M}w_{i}\cdot(y_{t,i}-y_{p,i})^{2},$ (8)
@@
$w_{i}=\begin{cases}10&\text{if }(y_{t,i}-y_{p,i})>0\\ 1&\text{if }(y_{t,i}-y_{p,i})\leq 0.\end{cases}$ (9)
This weighting approach introduces a bias for the overestimation of the safety function. Underestimation can lead to an orbit escaping early the region $Q$. Since the safety function defines the minimum control bound needed to confine the orbit, underestimations pose a critical risk. To mitigate this, the training process imposes a significantly higher penalty on underestimations, prompting the model to prioritize larger predictions. This conservative choice for the estimation lead to a satisfactory trade-off in performance as we will show in the next section.
4 Results
As detailed in Sec 3, the transformer-based model illustrated in Fig. 2 was trained on different amounts of sampled data. We trained $6$ models, each with progressively more data than the last. This approach simulates real world scenarios where data availability may be limited.